# Market Making without Regret

### v1

@@ Abstract
We consider a sequential decision-making setting where, at every round $t$, a market maker posts a bid price $B_{t}$ and an ask price $A_{t}$ to an incoming trader (the taker) with a private valuation for one unit of some asset. If the trader’s valuation is lower than the bid price, or higher than the ask price, then a trade (sell or buy) occurs. If a trade happens at round $t$, then letting $M_{t}$ be the market price (observed only at the end of round $t$), the maker’s utility is $M_{t}-B_{t}$ if the maker bought the asset, and $A_{t}-M_{t}$ if they sold it. We characterize the maker’s regret with respect to the best fixed choice of bid and ask pairs under a variety of assumptions (adversarial, i.i.d., and their variants) on the sequence of market prices and valuations. Our upper bound analysis unveils an intriguing connection relating market making to first-price auctions and dynamic pricing. Our main technical contribution is a lower bound for the i.i.d. case with Lipschitz distributions and independence between prices and valuations. The difficulty in the analysis stems from the unique structure of the reward and feedback functions, allowing an algorithm to acquire information by graduating the “cost of exploration” in an arbitrary way.
Keywords: Regret minimization, online learning, market making, first-price auctions, dynamic pricing.
@@ 1 Introduction
Trading in financial markets is a crucial activity that helps keep the world’s economy running, and several players, including hedge funds, prop trading firms, investment banks, central banks, and retail traders participate in it daily. While every actor has their own objective function (for example, a hedge fund wants to maximize profit whereas a central bank wants to keep inflation in check), at a fundamental level, trading can be viewed as a stochastic control problem where agents want the state to evolve so as to maximize their objective function. In this work we focus on market makers, i.e., traders whose job is to facilitate other trades to happen. One way to do so is by broadcasting, at all times, a price (bid) at which they are willing to buy, and a price (ask) at which they are willing to sell the asset being traded. This way when a buyer (seller) arrives, they do not have to wait for a seller (buyer) to be able to perform a transaction. A market where it is easy to make trades is called liquid, and liquidity is a desirable property for any kind of market. A market maker thus provides an essential service by increasing the liquidity of the market, and collects compensation for it by (among other ways) ensuring that the bid is always smaller than the ask, thus making a profit proportional to the difference between the bid and the ask. Market making is challenging, and a lot of thought goes into making it profitable, see [38] for an overview. A major risk that a market maker has to deal with, called adverse selection, is the risk that your counterpart is an informed trader who knows something about the future direction of the price movement. For example, an informed trader who knows that an asset is soon about to become cheaper will sell it to you thus forcing you to buy something whose price crashes. One way to mitigate this risk is to immediately offload your positions elsewhere. This strategy can be profitable if, for example, you have access to two markets for the same asset where one market is less liquid than the other. You can be a market maker for the less liquid market while offloading your positions in the more liquid one. We call this strategy “market making with instant clearing”.
@@ 1.1 Related works
Perhaps closest to our work is the paper [10], where they compete against a class of $N$ constant-spread dynamic strategies. Although we also compete with constant-spread (static) strategies, their results are not directly comparable to ours for several reasons. First, in their model the only unknown parameter is the market price, whose change from step to step is adversarial yet bounded by a known quantity. As the market price is revealed at the end of each step, their utility function at time $t$ is fully known at the end of step $t$. This implies that their model has a full information feedback, which yields a regret of order $\sqrt{T\ln N}$ against the best constant-spread strategy. In our setting, instead, we compete against a continuum of strategies parametrized by ask-bid pairs. This forces us to simultaneously control estimation and approximation error. Moreover, and crucially, our market model does not have a full information feedback, as our reward function depends on the taker’s private valuations which remain unknown. Hence, we need to carefully exploit the structure of the reward function to compensate for the missing information. Please see Section 5 for an extensive list of related works.
@@ 1.2 Our contributions
Adversarial IID
General Lip General Lip IV Lip+IV
Realistic $T$ $T^{\nicefrac{{2}}{{3}}}$ $T$ $T^{\nicefrac{{2}}{{3}}}$ $T^{\nicefrac{{2}}{{3}}}$ $T^{\nicefrac{{2}}{{3}}}$
Full $T$ $\sqrt{T}$ $\sqrt{T}$ $\sqrt{T}$ $\sqrt{T}$ $\sqrt{T}$
Table 1: A summary of the regret guarantees we prove for each variant of the online market-making problem. All the rates are optimal up to logarithmic factors.
We study the question of market making with instant clearing in an online learning setup. Here, trading happens in discrete time steps and an unknown stochastic process governs the prices of the asset being traded as well as the private valuations assigned to the asset by market participants (see details in Section 2). The market maker is an online learning algorithm that posts a bid and an ask at the beginning of each time step and receives some feedback at the end of the time step. We are interested in controlling the regret suffered by this learner at the end of $T$ time steps. We consider two feedback models. In both models, the learner can see the price (called the market price) at which they are able to offload their position at the end of the time step. However, in the realistic feedback model, they can only see whether their bid or the ask (or neither) got successfully traded, whereas in the full feedback model, they also get to see the private valuations of their counterpart in the trade. For each feedback model, we consider the i.i.d. setting, where the unknown stochastic process is i.i.d., as well as the adversarial setting, where no such assumption is made on the process. Our results are summarized in Table 1.
In the realistic feedback model, we make the following contributions.
@@
    We design the M3 (meta-)algorithm (Algorithm 2) and prove a $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$ upper bound on its regret under the assumption that either the cumulative distribution functions of the takers’ valuations are Lipschitz or the sequence of market prices and takers’ valuations are i.i.d. with market prices being independent of taker’s valuations (Theorem 3.3).
@@
    Our main technical contribution is a lower bound of order $\Omega(T^{\nicefrac{{2}}{{3}}})$ that matches M3’s upper bound, and holds even under the simultaneous assumptions that the sequence of market prices and takers’ valuations is i.i.d., admits Lipschitz cumulative distribution functions, and is such that market prices are independent of takers’ valuations (Theorem 3.4).
@@
    We then investigate the necessity of assuming either Lipschitzness of the cumulative distribution function of the takers’ valuations or independence of market values and takers’ valuations. We prove that, if both assumptions are dropped, learning is impossible in general, even when market values and takers’ valuations are i.i.d. (Theorem 3.5).
Lastly, we discuss the full-feedback case to flesh out the impact of limited feedback on learning rates and learnability. We show that learning is impossible in the adversarial setting (Theorem 4.1), while it is possible to achieve an $\mathcal{O}(\sqrt{T})$ regret rate when market values and takers’ valuations are i.i.d. (Theorem 4.3), or when the cumulative distribution functions of takers’ valuations are Lipschitz (Theorem 4.4). This rate is unimprovable, even if the three previous assumptions hold at the same time (Theorem 4.2).
1.3 Techniques and challenges
We now move on to describe the main technical challenges encountered when proving our results.
Upper bounds. First observe that the learner’s action space (i.e., the set of ask/bid pairs) is the two dimensional set $\{(b,a)\in[0,1]^{2}\mid b\leq a\}$. If market prices and takers’ valuations are both generated adversarially, standard bandit approaches typically require Lipschitzness (or the weaker one-side Lipschitzness) of the reward function, a property that is unfortunately missing in our setting. We can recover Lipschitzness (in expectation) by smoothing [^1] I.e., by assuming that their cumulative distribution function is Lipschitz. the taker’s valuations, while the market prices remain adversarial. Under the smoothness assumption, the problem can be seen as an instance of $2$-dimensional Lipschitz bandits, and a black-box application of existing continuous bandits techniques [60] would give us only a suboptimal regret rate $\mathcal{O}(T^{\nicefrac{{3}}{{4}}})$. Using a different approach that exploits the structure of our feedback and reward functions, we manage to effectively reduce the dimensionality of the problem by $1$, thus obtaining an improved (and optimal, as discussed later) $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$ regret rate, see Theorem 3.2.
The previous result is based on discretizing the action space and exploiting the Lipschitzness of the expected reward function provided by the smoothness assumption, together with the available feedback (which is sufficient to reconstruct bandit feedback). Can discretization be used to obtain the same rate when smoothed adversarial valuations are replaced by a different, yet natural, assumption? In Section 3.1 we consider an i.i.d. setting with independence between market prices and private valuations. Although this assumption does not provide Lipschitzness (not even in expectation), we show that the same algorithm as before has a regret again bounded by $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$. However, the analysis relies on a different observation: although we cannot guarantee that the reward of all actions is well approximated by that of a corresponding point in the grid, we can guarantee the weaker property that the expected reward of the best actions are approximated by the expected reward of a point in the grid, which is sufficient to prove the desired rate.
Lower bounds. A first roadblock in constructing regret lower bounds is that, in contrast to standard multi-armed bandits problems, we can control the distribution of the feedback and the expected reward functions only indirectly, through the joint distribution of market prices and valuations. In particular, there are constraints on the types of reward functions we can design, and it is non-trivial to design underlying joint distributions of market prices and valuations whose corresponding expected reward functions have the desired properties.
Another challenge is that the learner’s feedback in our setting is richer than bandit feedback. Therefore, even if we could encode our expected reward functions in hard multi-armed bandit instances, we would not be able to rely on the same arguments. A similar reason prevents us from applying the known $\Omega(T^{\nicefrac{{2}}{{3}}})$ lower bounds for dynamic pricing and first-price auctions, two problems that turn out to be closely related to ours. Proving the optimality of the $T^{\nicefrac{{2}}{{3}}}$ rate in the i.i.d. setting under the assumptions of smoothness and independence of market prices and taker’s valuations is a highly nontrivial task which represents the main technical contribution of this work.
To understand the complexity of proving a lower bound, we now compare the trade-offs between regret and amount of feedback in hard instances of our setting and of other related settings. The three main ingredients in the standard lower bound approach for online learning with partial feedback are:
  • The amount of information acquired when playing an action, quantified by the Kullback-Leibler (KL) divergence between the distribution of the feedback obtained when selecting the action in the perturbed instance and the corresponding feedback distribution when selecting the action in the base instance—the higher, the better.
  • The regret of an action, simply measured by the expected instantaneous regret of the action in the perturbed instance—the smaller, the better.
In hard instances of $K$-armed bandits, an arm is drawn uniformly at random from a $K$-sized set and assigned an expected reward $\varepsilon$-higher than the other arms in the set. Selecting an optimal arm costs zero regret and provides $\Theta(\varepsilon^{2})$ bits of information. Selecting a non-optimal arm, instead, provides no information and costs $\varepsilon$ regret [6]. This implies that any algorithm has only one way to acquire information about an arm, i.e., to play that arm. A similar situation occurs in other settings, including those with continuous decision spaces.
Figure 1: Utility (in red) and amount of acquired information (in blue) in a hard instance of the classic multi-armed bandit problem with a finite set of arms (on the left) and in a hard instance of the first-price auctions problem with a continuous set of bids (on the right). Both quantities are maximized at the same point. In particular where there is no perturbation, there is no information and some amount of reward, otherwise the KL can grow up to a quantity of order $\varepsilon^{2}$ and the utility can grow by a quantity of order $\varepsilon$.
For example, to acquire information about a bid in the standard hard instance of first-price auctions, the only reasonable option is to post that price, incurring regret $\Theta(\varepsilon)$ if the arm is non-optimal and acquiring $\Theta(\varepsilon^{2})$ bits of information if the arm is optimal [25]. See Figure 1 for a pictorial representation of the expected reward and KL divergence in said hard instances.
In a hard instance of market-making, a price $b^{\star}$ is drawn uniformly at random from a known finite set of bid prices and the bid/ask pair $(b^{\star},1)$ is given a reward $\varepsilon$-higher than the other bid/ask pairs of the form $(b^{\prime},1)$. The amount of information and corresponding regret of each pair of prices $(b,a)$ can be visualized by looking at the qualitative plots in Figure 3. Note that there are uncountably many ways to determine if a pair of prices $(b,1)$ is optimal: playing any pair in the non-blue area of the left plot in Figure 3 allows to modulate the amount of exploration, with the points that yield the highest quality feedback being the ones with low reward and high regret (for a plot representing the growth of the amount of feedback and the corresponding reward as a function of the pair $(b,a)$ taken along a representative path, see Figure 4). Consequently, the exploration-exploitation trade-off consists in first choosing whether to exploit with some pair $(b,1)$, or instead pick another potentially optimal pair $(b^{\prime},1)$ to explore, then choosing from an uncountable set of options $(b^{\prime\prime},a^{\prime\prime})$ how much regret one is willing to suffer in order to acquire more information about the optimality of $(b^{\prime},1)$.222 The situation is even more complex because, in the realistic feedback model, pairs $(b,a)$ allow to test for the optimality not just of the pair $(b,1)$, but also of the pair $(a,1)$. To build a lower bound, we then need to consider all possible ways a learner can handle these multiple trade-offs between information and regret. To the best of our knowledge, this is the first lower bound for a problem exhibiting a continuum of trade-offs between reward and feedback. As our setting is arguably simple, we expect the same phenomenon to occur in other applications of online learning to digital markets. We solve the problem by partitioning the action space into finitely many elements (or clusters) that can be further grouped in three macro-categories: “get high regret but (potentially) lots of information”, “get (potentially) small regret but also little information”, and clearly suboptimal actions where we “get some regret and no information”. We then analyze an algorithm in terms of how many times it selects an action from each cluster, using similar techniques for all clusters belonging to the same macro-category. We believe that the clustering technique developed in this work could be helpful in tackling similar lower bounds exhibiting a continuum of trade-offs that might appear in future works applying online learning to digital market problems.
We close this section with a last remark. When investigating the necessity of the assumptions of smoothness and independence of valuations and prices, if both these assumptions are dropped, an intriguing needle-in-a-haystack phenomenon arises: by devising discrete entangled takers’ valuations and market prices distributions, we show that it becomes impossible for the learner to determine an optimal action within a finite horizon in a continuum of potentially optimal actions.
@@ 2 Setting
In online market making, the action space of the maker is the upper triangle $\mathcal{U}\coloneqq\bigl{\{}(b,a)\in[0,1]^{2}\,:\,b\leq a\bigr{\}}$, enforcing the constraint that a bid price $b$ is never larger than the corresponding ask price $a$. For an overview of the notation, see Table 2, in Appendix A. The utility of the market maker, for all $(b,a)\in\mathcal{U}$ and $m,v\in[0,1]$, is
$u(b,a,m,v)\coloneqq(m-b)\cdot\underbrace{\mathbb{I}\{b\geq v\}}_{\text{Maker buys}}+\;(a-m)\cdot\underbrace{\mathbb{I}\{a<v\}}_{\text{Maker sells}}\;,$ (1)
where $v$ is the taker’s private valuation and $m$ is the market price. [^3] The choice of buying when $b\geq v$ and selling when $a<v$ models a taker that is slightly more inclined to sell rather than buy: in this case, the taker is willing to sell even when their valuation is exactly equal to the bid. This choice is completely immaterial for the results that follow and is merely done for the sake of simplicity. All the results we present still hold if trades happen according to a similar rule but with a taker that is slightly more inclined to buy rather than sell (i.e., if a trade occurs whenever $b>v$ or $a\leq v$) or if this inclination changes arbitrarily whenever a new taker comes at any new time step.
for time $t=1,2,\ldots$  do
      Taker arrives with a private valuation $V_{t}\in[0,1]$;
      Maker posts bid/ask prices $(B_{t},A_{t})\in\mathcal{U}$;
      Feedback $\mathbb{I}\{B_{t}\geq V_{t}\}$ and $\mathbb{I}\{A_{t}<V_{t}\}$ is revealed;
      if $B_{t}\geq V_{t}$ then Maker buys and pays $B_{t}$ to Taker;
      else if $A_{t}<V_{t}$ then Maker sells (short) and is paid $A_{t}$ by Taker;
      else no trade happens;
      The market price $M_{t}\in[0,1]$ is revealed;
     if Maker bought from Taker then Maker offloads to Market at price $M_{t}$;
     else if Maker sold (short) to Taker then Maker buys from Market at price $M_{t}$;
      Maker’s utility is $U_{t}(B_{t},A_{t})\coloneqq u(B_{t},A_{t},M_{t},V_{t})$;
tradingprotocol]a:trading-protocol
Trading protocol 1 Market making with instant clearing and realistic feedback
Our online market making protocol is specified in LABEL:a:trading-protocol. At every round $t$, a taker arrives with a private valuation $V_{t}\in[0,1]$ and the maker posts bid/ask prices $(B_{t},A_{t})\in\mathcal{U}$. If $B_{t}<V_{t}\leq A_{t}$ (i.e., if the taker is not willing to sell nor to buy at the proposed bid/ask prices), then no trade happens. If $B_{t}\geq V_{t}$ (buy) or $A_{t}<V_{t}$ (sell), a trade happens. At the end of each round, the maker observes the market price $M_{t}$ and the type of trade (buy, sell, none) that took place in that round. The maker’s utility is determined by $U_{t}(B_{t},A_{t})\coloneqq u(B_{t},A_{t},M_{t},V_{t})$. Hence, any (possibly randomized) learning strategy for the maker at time $t$ is a map that takes all previous feedback $\big{\{}\big{(}\mathbb{I}\{B_{\tau}\geq V_{\tau}\},\mathbb{I}\{A_{\tau}<V_{\tau}\},M_{\tau}\big{)}\big{\}}_{\tau<t}$, plus some optional random seeds, to a (random) pair $(B_{t},A_{t})$ of bid/ask prices in $\mathcal{U}$.
@@
where the expectations are with respect to the (possible) randomness of the market prices, takers’ valuations, and (possibly) the internal randomization of the algorithm. Note that the supremum is not attained in general, because of the strict inequality in one of the two indicator functions in Equation 1.
Types of feedback.
We call the feedback received in the protocol above “realistic feedback”. In a real market, we very rarely have access to the private valuation $V_{t}$ of the taker and thus it is safe to assume that the learner only gets to observe $\mathbb{I}\{B_{t}\geq V_{t}\}$ and $\mathbb{I}\{A_{t}<V_{t}\}$. Still, one can imagine scenarios where additional information about $V_{t}$ is sometimes available. In this spirit and with the goal of contrasting the effect of partial feedback on learnability and learning rates, it is of theoretical interest to study what happens in the other extreme, the full feedback scenario, where the learner gets to observe the value of $V_{t}$ (in addition to $M_{t}$) at the end of each time step $t$. There might be other interesting feedback models between realistic and full that we leave for future work.
3 Realistic feedback

### v2

@@ Abstract
We consider a sequential decision-making setting where, at every round $t$, the learner (a market maker) posts a bid price $B_{t}$ and an ask price $A_{t}$ to an incoming trader (the taker) with a private valuation for some asset. If the trader’s valuation is lower than the bid price, or higher than the ask price, then a trade (sell or buy) occurs. Letting $M_{t}$ be the market price (observed only at the end of round $t$), the maker’s utility is $M_{t}-B_{t}$ if the maker bought the asset, it is $A_{t}-M_{t}$ if they sold it, and it is 00 if no trade occurred. We characterize the maker’s regret with respect to the best fixed choice of bid and ask pairs under a variety of assumptions (adversarial, i.i.d., and their variants) on the sequence of market prices and valuations. Our upper bound analysis unveils an intriguing connection relating market making to first-price auctions and dynamic pricing. Our main technical contribution is a lower bound for the i.i.d. case with Lipschitz distributions and independence between market prices and takers’ valuations. The difficulty in the analysis stems from a unique relationship between the reward and feedback functions that allows learning algorithms to trade off reward for information in a continuous way.
keywords:
Regret minimization, online learning, market making, first-price auctions, dynamic pricing.
@@ 1 Introduction
Trading in financial markets is a crucial activity that helps keep the world’s economy running, and several players, including hedge funds, prop trading firms, investment banks, central banks, and retail traders participate in it daily. While every actor has their own objective function (for example, a hedge fund wants to maximize profit whereas a central bank wants to keep inflation in check), at a fundamental level, trading can be viewed as a stochastic control problem where agents want the state to evolve so as to maximize their objective function. In this work we focus on market makers, i.e., traders whose job is to facilitate other trades to happen. One way to do so is by broadcasting, at all times, a price (bid) at which they are willing to buy, and a price (ask) at which they are willing to sell the asset being traded. This way when a buyer (resp., seller) arrives, they do not have to wait for a seller (resp., buyer) to be able to perform a transaction. A market where it is easy to make trades is called liquid, and liquidity is a desirable property for any kind of market. A market maker thus provides an essential service by increasing the liquidity of the market. Market making is challenging, and a lot of thought goes into making it profitable, see Harris (2003) for an overview. A major risk that a market maker has to deal with, called adverse selection, is the risk that your counterpart is an informed trader who knows something about the future direction of the price movement. For example, an informed trader who knows that an asset is soon about to become cheaper will sell it to you thus forcing you to buy something whose price will crash. One way to mitigate this risk is to immediately offload your positions elsewhere. More generally, a market maker does not specialize in predicting the future movements of prices, and thus dislikes holding assets for too long. Market making desks often have limits on how much inventory they are allowed to hold at a time (Fender and Lewrick, 2015).
The strategy of immediately offloading your position, though simple, is profitable if one does it strategically. Consider, for example, the case where one can trade the same asset on two exchanges, where its bid-ask spread on the first is smaller than its bid-ask spread on the second. In this case, one can make a market on the second while offloading their positions on the first. This situation arises frequently for two closely related assets. For example, a futures contract on an asset is often more liquid than the asset itself and their price movements are highly correlated. For these reasons, we focus on a setting where the market maker offloads their positions shortly after each transaction.
@@ 1.1 Related works
Online market making.
Our work is closely related to that of Abernethy and Kale (2013), which analyzes competition against $N$ constant-spread dynamic strategies. While we also consider constant-spread (static) strategies, key differences make their results non-comparable to ours. Their model assumes full information feedback, with an adversarial yet bounded market price change and a fully known utility function at each step, leading to a regret of $\sqrt{T\ln N}$. In contrast, we compete against a continuum of ask-bid strategies while managing both estimation and approximation errors. Crucially, our model lacks full information feedback, as rewards depend on private valuations, requiring structural exploitation of the reward function to mitigate uncertainty.
Our work is also related to the classic Glosten-Milgrom model (Glosten and Milgrom, 1985), which introduced the concept of a market specialist (maker) providing liquidity to the market within an exchange. The specialist interacts with both informed traders (whose valuations are influenced by future market prices) and uninformed ones. The specialist’s objective is to distinguish between informed and uninformed traders via interactions in an online fashion and optimize the bid-ask spread accordingly. In 2005, Das considers a learning algorithm in an extended version of the Glosten-Milgrom model for the market maker with a third type of “noisy-informed” trader, whose current valuation is a noisy version of the future market price. However, to the best of our knowledge, they do not provide any regret guarantees.
Smoothed adversary.
Popularized by Spielman and Teng (2004); Rakhlin et al. (2011); Haghtalab et al. (2020), smoothness analysis provides a framework for analyzing algorithms in problems parameterized by distributions that are not “too” concentrated. Recent advancements in the smoothed analysis of online learning algorithms include contributions by Kannan et al. (2018); Haghtalab et al. (2022, 2021); Block et al. (2022); Durvasula et al. (2023); Cesa-Bianchi et al. (2021, 2023); Bolić et al. (2024); Cesa-Bianchi et al. (2024a). In this work, we exploit the connection between the smoothness of the takers’ valuations distributions (i.e., the Lipschitzness of the corresponding cumulative distribution function, or, equivalently, the boundedness of the corresponding probability density function) and Lipschitzness of the expected utility. This property is crucial to achieving sublinear regret guarantees in adversarial settings and affects learning depending on different assumptions on the level of information shared with the learner while bidding.
Please see Appendix B for further related works.
Online learning in digital markets.
Our work unveils a connection between market making and first-price auctions/dynamic pricing, thus contributing to the existing body of research on the multi-armed bandit framework, which boasts a substantial literature base across statistics, operations research, computer science, and economics (Auer et al., 2002a; Lattimore and Szepesvári, 2020; Badanidiyuru et al., 2018; Kleinberg, 2004). The multi-armed bandit framework has proven to be an effective tool for addressing key problems in finance under uncertainty, including repeated first-price auctions (Cesa-Bianchi et al., 2024c; Han et al., 2024; Achddou et al., 2021; Flajolet and Jaillet, 2017) and dynamic pricing (Cesa-Bianchi et al., 2019; Babaioff et al., 2015; Kleinberg and Leighton, 2003). These settings are closely related to online market making (see Appendix B for further discussion).
Closely related to dynamic pricing is the problem of optimal taxation (Cesa‐Bianchi et al., 2025), where the interaction between the learner and the environment follows a similar structure, but the objective shifts from revenue maximization to the design of pricing mechanisms that maximize social welfare.
Beyond first-price auction settings, related auction formats include repeated Vickrey auctions, studied from both the bidder’s perspective (Weed et al., 2015) and the seller’s perspective (Cesa-Bianchi et al., 2013).
Another important problem at the intersection of online learning and digital markets is repeated bilateral trade. In this setting, the learner acts as a broker, proposing price pairs to a sequence of traders with unknown private valuations. The foundational works of Cesa-Bianchi et al. (2021, 2024b) established optimal regret rates under i.i.d. valuations, assuming smoothness and independence between buyer and seller values, using the gain from trade as reward function, and mechanisms that are strongly budget balanced. This line of work has been extended in multiple directions: Azar et al. (2022) proposed weakly budget-balanced mechanisms achieving optimal $2$-regret in adversarial settings; Cesa-Bianchi et al. (2023, 2024a) relaxed the i.i.d. and the independence between sellers and buyers assumptions while maintaining optimal regret guarantees under smoothness alone; and Bernasconi et al. (2024) introduced globally budget-balanced mechanisms with sublinear regret even in adversarial settings.
Contextual versions of bilateral trade have also been studied by Gaucher et al. (2025), while Bachoc et al. (2024) took a different perspective by proposing a fairness-oriented reward function establishing optimal regret bounds under the i.i.d. assumption with independent buyer and seller valuations, using strongly budget-balanced mechanisms.
Recent work has also considered scenarios in which agents are not assigned fixed roles as buyers or sellers. In this more general setting, Bolić et al. (2024) achieved faster regret rates than in standard bilateral trade using strongly budget-balanced mechanisms under the same assumptions as Cesa-Bianchi et al. (2021). In the contextual setting, Bachoc et al. (2025b) obtained optimal fast rates under a noisy-linear model, and Bachoc et al. (2025a) extended these results to nonparametric contexts. Finally, Cesari and Colomboni (2025) studied the objective of maximizing the volume of trade, achieving accelerated regret rates in this setting.
@@ 1.2 Our contributions
We study the question of market making in an online learning setup. Here, trading happens in discrete time steps and an unknown stochastic process governs the prices of the asset being traded as well as the private valuations assigned to the asset by market participants (see details in Section 2). The market maker is an online learning algorithm that posts a bid and an ask at the beginning of each time step and receives some feedback at the end of the time step. We are interested in controlling the regret suffered by this learner at the end of $T$ time steps. At the end of each round, the learner observes the market price, i.e., the price at which they are able to offload their position. Additionally, they see whether their bid or their ask (or neither) got successfully traded. We consider settings where the unknown stochastic process is independent and identically distributed (iid), as well as the adversarial setting (adv), where no such assumption is made on the process. We also consider settings where the cumulative distribution functions of takers’ valuations are Lipschitz (lip), as well as settings where at each time step $t$ the corresponding market price and taker’s valuation are independent of each other (iv). Our results are summarized in Table 1.
Setting (adv) (adv+lip) (iid) (iid+lip) (iid+iv) (iid+lip+iv)
Regret $T$ $T^{\nicefrac{{2}}{{3}}}$ $T$ $T^{\nicefrac{{2}}{{3}}}$ $T^{\nicefrac{{2}}{{3}}}$ $T^{\nicefrac{{2}}{{3}}}$
Table 1: A summary of the regret guarantees we prove for each variant of the online market-making problem. All the rates are optimal up to constant factors.
We make the following contributions.
@@
    We design the M3 (meta-)algorithm (Section 3) and prove a $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$ upper bound on its regret under the assumption that either the cumulative distribution functions of the takers’ valuations are Lipschitz (lip) or the sequence of market prices and takers’ valuations are independent and identically distributed (iid) with market prices being independent of takers’ valuations (iv) (Theorem 3.1).
@@
    Our main technical contribution is a lower bound of order $\Omega(T^{\nicefrac{{2}}{{3}}})$ that matches M3’s upper bound, and holds even under the simultaneous assumptions that the sequence of market prices and takers’ valuations is independent and identically distributed (iid), market prices are independent of takers’ valuations (iv), the cumulative distribution function of the takers’ valuations is Lipschitz (lip), and the cumulative distribution function of the market prices is Lipschitz (Theorem 4.1).
@@
    We then investigate the necessity of assuming either Lipschitzness of the cumulative distribution function of the takers’ valuations (lip) or the independence of market prices and takers’ valuations (iv). We prove that, if both assumptions are dropped, learning is impossible in general, even when market prices and takers’ valuations are independent and identically distributed (iid) (Theorem 5.1).
We can interpret the independence between takers’ valuation and market prices (iv) as the taker being an uninformed trader (in the same spirit as the Glosten-Milgrom model). If we consider the (iid) setting, and contrast the impossibility of learning when neither Lipschitzness (lip) nor independence (iv) is assumed with the $\mathcal{O}(T^{\nicefrac{{2}}{{3}}})$ upper bound when the market prices are independent of takers’ valuations (iv), we get evidence for the practitioner’s intuition that the existence of uninformed traders is good for the market maker.
@@ 2 Setting
{algorithm2e}
[t]
[…]
$u(b,a,m,v)\coloneqq\underbrace{(m-b)}_{\text{market-bid spread}}\!\!\!\!\times\;\underbrace{\mathbb{I}\{b\geq v\}}_{\text{Maker buys}}\;\;\;+\underbrace{(a-m)}_{\text{ask-market spread}}\!\!\!\!\times\;\underbrace{\mathbb{I}\{a<v\}}_{\text{Maker sells}}\;,$ (1)
where $v$ is the taker’s private valuation and $m$ is the market price. [^2] The choice of buying when $b\geq v$ and selling when $a<v$ models a taker that is slightly more inclined to sell rather than buy: in this case, the taker is willing to sell even when their valuation is exactly equal to the bid. This choice is completely immaterial for the results that follow and is merely done for the sake of simplicity. All the results we present still hold if trades happen according to a similar rule but with a taker that is slightly more inclined to buy rather than sell (i.e., if a trade occurs whenever $b>v$ or $a\leq v$) or if this inclination changes arbitrarily whenever a new taker comes at any new time step. Our online market making protocol is specified in Section 2. At every round $t$, a taker arrives with a private valuation $V_{t}\in[0,1]$ and the maker posts bid/ask prices $(B_{t},A_{t})\in\mathcal{U}$. If $B_{t}<V_{t}\leq A_{t}$ (i.e., if the taker is not willing to sell nor to buy at the proposed bid/ask prices), then no trade happens. If $B_{t}\geq V_{t}$ (buy) or $A_{t}<V_{t}$ (sell), a trade happens. At the end of each round, the maker observes the market price $M_{t}$ and the type of trade (buy, sell, none) that took place in that round. For each $(b,a)\in\mathcal{U}$ and each round $t$, defining $U_{t}(b,a)\coloneqq u(b,a,M_{t},V_{t})$, the maker’s utility at round $t$ is $U_{t}(B_{t},A_{t})=u(B_{t},A_{t},M_{t},V_{t})$. Hence, any (possibly randomized) learning strategy for the maker at time $t$ is a map that takes as input past feedback $\big{\{}\big{(}\mathbb{I}\{B_{\tau}\geq V_{\tau}\},\mathbb{I}\{A_{\tau}<V_{\tau}\},M_{\tau}\big{)}\big{\}}_{\tau<t}$, plus some optional random seeds, and returns as output a (random) pair $(B_{t},A_{t})$ of bid/ask prices in $\mathcal{U}$.
@@
where the expectations are with respect to the (possible) randomness of the market prices, takers’ valuations, and (possibly) the internal randomization of the algorithm.
We study this problem under a mix of several standard assumptions on the underlying process $(V_{t},M_{t})_{t\in\mathbb{N}}$ modeling private valuations and market prices.
  • (adv)
    $(V_{t},M_{t})_{t\in\mathbb{N}}$ is a stochastic process.
  • (iid)
    $(V_{t},M_{t})_{t\in\mathbb{N}}$ is an i.i.d. stochastic process. [^3] Note that this assumption alone does not require that, for any $t\in\mathbb{N}$, $V_{t}$ is independent of $M_{t}$.
  • (iv)
    $(V_{t},M_{t})_{t\in\mathbb{N}}$ is a stochastic process such that, for all $t\in\mathbb{N}$, $V_{t}$ is independent of $M_{t}$. [^4] Note that this assumption alone does not require that $(V_{t},M_{t})_{t\in\mathbb{N}}$ is an independent stochastic process.
  • (lip)
    $(V_{t},M_{t})_{t\in\mathbb{N}}$ is a stochastic process such that, there exists $L>0$ such that, for all $t\in\mathbb{N}$, the cumulative distribution function of $V_{t}$ is $L$-Lipschitz.
The variables $A_{t},B_{t},V_{t},M_{t}$ can be thought of as relative to a global reference value revealed at the beginning of each time step, aligning with the convention of quoting in “points” (or “ticks”) above or below a benchmark. With this interpretation, in the regret definition, the benchmark strategy maintains a constant distance from a reference throughout the trading period (as opposed to a constant absolute level). These observations greatly broaden the scope of application of our setting.
[…]
Since the learner’s action space contains a continuum of actions, standard bandit approaches require Lipschitzness (or the weaker one-side Lipschitzness) of the reward function, a property that is unfortunately missing in our setting, even when the environment is stable (iid) and the takers are uninformed (iv). One way to recover Lipschitzness (although only in expectation) is by assuming the takers’ valuations are smoooth, i.e., by assuming that their cumulative distribution functions are Lipschitz (lip). Under the (lip) assumption, recalling that our action space $\mathcal{U}$ is the set of bid/ask pairs $\{(b,a)\in[0,1]^{2}\mid b\leq a\}$, and noticing that the feedback the learner receives at the end of each time step is sufficient to reconstruct bandit feedback, the market making problem can be seen as an instance of $2$-dimensional Lipschitz bandits. However, a black-box application of existing $d$-dimensional Lipschitz bandits techniques (Slivkins, 2019, with $d=2$) would give only a suboptimal regret rate of $\mathcal{O}(T^{\nicefrac{{(d+1)}}{{(d+2)}}})=\mathcal{O}(T^{\nicefrac{{3}}{{4}}})$.